OpenStudy (anonymous):

For f(x) = e^sin(x) use your graphing calculator to find thr number of zeros of f'(x) on the clised interval [0,2π]

OpenStudy (zale101):

f(x) = e^sin(x), take the derivative (by applying the chain rule), set it equal to zero. See how many solutions for x you'll get. @familyguymath

OpenStudy (anonymous):

\(\large f^\prime (x) = \frac{ d }{ dx } (e^u) . \frac{ d }{ dx } (u)\) whic \(( u=sinx) \)

OpenStudy (anonymous):

u can also use the graph of the function as the question wanted u ;)

OpenStudy (anonymous):

@familyguymath